# fourierlargesymbols147 thanks: Present address for Q. H. Tran: Quantum Laboratory, Fujitsu Research, Fujitsu Limited

### v4

@@ III Conclusion
This paper proposes non-stationary and subspace/subset ESPs that are helpful in real-world RC scenarios. As a concrete application, we numerically analyzed a QRC with a well-known SK Hamiltonian, a reset-input encoding method for non-stationary ESP, and an adaptive entanglement system for a subset of non-stationary ESP. Our study revealed a partial correspondence between the non-stationary ESP and the subset/subspace ESP and the information processing capabilities, as demonstrated using NARMA tasks and MC/IPC calculations.
@@ Appendix A Equivalence of ESP definitions
Here, we prove the equivalence of two distinct ESP definitions introduced in [4] so that our following discussion becomes general.
@@
    Let a state sequence given inputs $\{\mathbf{u}_{t}\}$ be $\{\mathbf{x}_{t}\}$ such that $\mathbf{x}_{t}=T(\mathbf{x}_{t-1},\mathbf{u}_{t})$. Let $\{\mathbf{x}_{t}\}$ and $\{\mathbf{x}_{t}^{\prime}\}$ be two different state sequences with the same input sequence and different initial states. Then,

### v5

@@ III Conclusion
This paper proposes non-stationary, subspace, and subset ESPs that are helpful in real-world RC scenarios. As a concrete application, we numerically analyzed a QRC with a well-known SK Hamiltonian, a reset-input encoding method for non-stationary ESP, and an adaptive entanglement system for a subset of non-stationary ESP. Our study revealed a partial correspondence between the non-stationary ESP, the subset and subspace ESP, and the information processing capabilities, as demonstrated using NARMA tasks and MC/IPC calculations.
@@ Appendix A Equivalence of ESP definitions
Here, we prove the equivalence of two distinct ESP definitions introduced in [4] so that our discussion becomes general.
@@
    Let a state sequence given inputs $\{\mathbf{u}_{t}\}$ be $\{\mathbf{x}_{t}\}$ such that $\mathbf{x}_{t}=T(\mathbf{x}_{t-1},\mathbf{u}_{t})$ for some $T:\mathcal{X}\times U\to\mathcal{X}$. Let $\{\mathbf{x}_{t}\}$ and $\{\mathbf{x}_{t}^{\prime}\}$ be two different state sequences with the same input sequence and different initial states. Then,
@@
Appendix B ESP results for the other random Hamiltonians
We calculated the ESP and non-stationary ESP indicators for QRCs with SK Hamiltonians with different parameters from the one we analyzed in Sec. II.3. Each Hamiltonian parameter configuration is represented by: $H_{2}(h\sim 0.013$ and $W\sim 1.05)$, $H_{3}(h\sim 0.377$ and $W\sim 24.8)$, $H_{4}(h\sim 57.2$ and $W\sim 47.5)$ and $H_{5}(h\sim 48.3$ and $W\sim 0.0305)$.
As shown in Fig. 14 below, after feeding the appropriate length of input signals, every setup reaches the point where the upper and bottom line is the only region in which the non-stationary ESP is considerably weaker compared to with other regions. The length input sequence required to reach that point is governed by the respective decay strength of fading memory in each Hamiltonian: for instance, $H_{3}$ and $H_{4}$ relate to the many-body localization phase as shown in [32], so it requires a longer input sequence to reach that point.
((a))
((b))
((c))
((d))
((e))
((f))
((g))
((h))
((i))
((j))
((k))
((l))
Figure 14: Non-stationary ESP indicators for QRCs driven by the random Hamiltonians $H_{2}$ (a-c), $H_{3}$ (d-f), $H_{4}$ (g-i), and $H_{5}$ (j-l). $t=200$ for the left column (a, d, g, j), $t=2000$ for the middle column (b, e, h, k), and $t=20000$ for the right column (c, f, i, l), where $t$ represents the time steps used in the calculation of the non-stationary ESP indicator (Eq. 18).